# DAOcc: 3D Object Detection Assisted Multi-Sensor Fusion for 3D Occupancy Prediction

### v1

@@ 3.3 BVRE
Considering the sparsity of point clouds, the increase in computational cost resulting from expanding the processing range of point clouds is very small compared to images. Therefore, we extend the point cloud range to provide more 3D spatial context, compensating for the information loss caused by lower image resolution. However, arbitrarily setting the $XY$ range of the point cloud or voxel resolution may result in misalignment between the BEV features and the occupancy ground truth. Assuming that the voxel resolution of the occupancy grid is $res_{o}$, the expanded point cloud data in BEV covers a rectangular region with $x$-coordinates ranging from $-x$ to $x$ and $y$-coordinates ranging from $-y$ to $y$, and has a voxel resolution of $res_{p}$. Thus, $x$ and $y$ must be integer multiples of $res_{o}$, and $res_{o}$ must also be an integer multiple of $res_{p}$. Otherwise, the spatial resolution along the $z$-axis for each feature vector in the BEV representation will not be equal to $res_{o}$. To avoid complex manual design, we adopted coordinate transformation and interpolation, as illustrated in the purple part of Figure 2, which can be formulated as follows:
$F_{occ}={\rm GridSample}(F_{r},{\rm Norm}(T_{o2p}\times P_{o}))$ (1)
where $P_{o}$ is a set of predefined points in the $XY$ plane of the coordinate system of the occupancy annotation, each point is located at the center of an occupancy grid in the $XY$ plane. $T_{o2p}$ is the transformation matrix from occupancy coordinates to Lidar coordinates. The function of ${\rm Norm}$ is to scale the coordinate values to a range from -1 to 1.
3.4 Detachable Auxiliary Head
To fully exploit the inherent geometry information in point clouds, we add a separable auxiliary detection network to the fused features, which will enable the features to benefit more from the geometry information. Meanwhile, the auxiliary detection task is related to the occupancy prediction task, thus providing multiple regularization effects during optimization. For simplicity, we use the one-stage CenterHead introduced in CenterPoint [53] as the auxiliary detection head. Given the refined BEV feature $F_{r}$, we utilize two convolutional layers to generate a keypoint heatmap $HM={p_{xy}}$ from the BEV perspective, and apply a Gaussian kernel [17, 60] to map the center points of all ground truth 3D boxes onto a target heatmap $T$. The training objective is formulated as a focal loss [38] based on Gaussian heatmaps:
@@ 3.4 Detachable Auxiliary Head
$\mathcal{L}_{det}=\mathcal{L}_{cls}+\lambda_{r}\mathcal{L}_{loc}$ (4)
@@ 3.5 Overall Objective Function
Most existing methods rely on complex combinations of loss functions [46, 47, 23, 55, 39, 50, 33, 58, 34, 40], such as a combination of several or all of the focal loss [38], the scene-class affinity loss [4], the dice loss, the lovasz-softmax loss [2], the depth loss [21], etc., to achieve the expected performance. In contrast, our approach, utilizing the proposed auxiliary supervision for 3D object detection, requires only a simple cross-entropy loss $\mathcal{L}_{ce}$ for occupancy prediction. The total loss for our framework can be defined as:
@@ 4 Experiments
Method mIoU Modality Input Size 2D Backbone
@@
FlashOcc(2f)* [54] 43.52 C 512 $\times$ 1408 Swin-B 13.4 51.1 27.7 51.6 56.2 27.3 30.0 29.9 29.8 37.8 43.5 83.8 46.6 56.2 59.6 50.8 44.7
OSP [39] 39.41 C 900 $\times$ 1600 R101-DCN 11.2 47.3 27.1 47.6 53.7 23.2 29.4 29.7 28.4 32.4 39.9 79.4 41.4 50.3 53.2 40.5 35.4
DHD(2f) [50] 45.53 C 512 $\times$ 1408 Swin-B 14.1 53.1 32.4 52.4 57.4 30.8 35.2 33.0 33.4 37.9 45.3 84.6 48.0 57.4 60.3 52.3 46.2
@@
DAOcc 53.82 C+L 256 $\times$ 704 R50 12.4 59.6 38.4 61.9 67.1 35.3 48.2 59.1 43.5 50.9 56.3 83.0 44.7 56.7 59.9 70.0 68.1
Table 1: 3D Semantic occupancy prediction performance on Occ3D-nuScenes validation set. (xf) means use x frames for temporal fusion. * denotes the usage of exponantial moving average (EMA). The C, L, and R denote camera, LiDAR, and radar, respectively.
Method RayIoU Modality Input Size 2D Backbone RayIoU1m RayIoU2m RayIoU4m mIoU
@@
DAOcc 48.2 C+L 256 $\times$ 704 R50 44.4 48.7 51.6 47.88
Table 2: 3D Semantic occupancy prediction performance on Occ3D-nuScenes validation set without camera mask. (xf) means use x frames for temporal fusion. The C and L denote camera, LiDAR, respectively.
Method IoU mIoU Modality Input Size 2D Backbone
@@
DAOcc 45.0 30.5 C+L 256$\times$ 704 R50 30.8 19.5 34.0 38.8 25.0 27.7 29.9 22.5 23.2 31.6 41.0 25.9 29.4 29.9 35.2 43.5
Table 3: 3D Semantic occupancy prediction performance on SurroundOcc validation set. The C, L, and R denote camera, LiDAR, and radar, respectively.
@@ 4.1 Datasets and Evaluation Metrics
Occ3D-nuScenes [41] is a large-scale benchmark for 3D occupancy prediction, containing 700 scenes for training and 150 for validation, each lasting 20 seconds with annotations at 2 Hz. The dataset covers a perception range of [-40m, 40m] in the $X$ and $Y$ directions and [-1m, 5.4m] along the $z$-axis, which is discretized into voxels of [0.4m, 0.4m, 0.4m] in size. Each occupied voxel is assigned one of 17 semantic labels, including 16 common classes and a general object class ‘others’. The data includes LiDAR point clouds and RGB images from six surround cameras, enabling dense voxel-wise annotation for 3D scene understanding. In addition, the dataset also provides visibility masks for LiDAR and camera modes, and it can be used for training.
SurroundOcc [48] is an automatically generated occupancy prediction dataset without human annotation. It is built upon nuScenes [3] and generates dense occupancy ground truth utilizing existing 3D detection and 3D semantic segmentation labels. The dataset employs the training set to train the model and utilizes the validation set for evaluation purposes. The occupancy prediction range is set to [-50m, 50m] for the $X$ and $Y$ axes, and [-5m, 3m] for the $z$-axis. The final output occupancy grid has a shape of 200×200×16, with a voxel size of [0.5m, 0.5m, 0.5m].
Evaluation metrics: Following the previous methods, when training Occ3D-nuScenes with the camera visible mask, we use the mean Intersection over Union (mIoU) for fair comparison with previous methods. When the camera visible mask is not used during training, we report an additional semantic segmentation metric, RayIoU [29], which was recently proposed. For SurroundOcc, we report semantic segmentation performance using IoU and mIoU, where IoU is the intersection over union of occupied voxels. This IoU is used as an evaluation metric for the scene completion (SC) task by ignoring the semantic categories of occupied voxels.
@@ 4.2 Implementation Details
We implement our method using PyTorch [36] and train all models on 8 NVIDIA RTX 4090 GPUs. We utilize ResNet-50 [11], which is pre-trained on nuImages [3], as the image backbone and crop the input image size to 256×704. In the LiDAR branch, we voxelize 10 LiDAR scans and start training from scratch. We employ the AdamW [32] optimizer with a cosine annealing learning rate scheduler, including a warmup phase, and set an initial learning rate of 2e-4. We apply data augmentation to both the input image and the ground truth 3D bounding boxes, and adopt CBGS [61].
Figure 3: Qualitative visualizations on Occ3D-nuScenes validation set without camera mask.
@@ 4.3 Comparison with State-of-the-Art Methods
Comparison on Occ3D-nuScenes with camera mask. As presented in Table 1, we compare our DAOcc with both single-modal and multi-modal methods on Occ3D-nuScenes, all results are either implemented directly by their authors or reported based on their official code implementations. All methods use the camera visible mask during training, except for CTF-Occ [41]. Our DAOcc establishes a new state-of-the-art performance, achieving an outstanding 53.82 mIoU, while utilizing a 2D backbone ResNet50 and 256×704 input image resolution. Compared with the current state-of-the-art multi-modal method EFFOcc [40], our DAOcc outperforms EFFOcc by 1.2 mIoU. Moreover, while EFFOcc employs Swin-B and an input image resolution of 512×1408, DAOcc utilizes only ResNet50 and an input image resolution of 256×704. In addition, EFFOcc requires the LiDAR branch to perform 3D object detection pre-training, which results in an improvement of approximately 3 mIoU, in order to achieve optimal results, whereas DAOcc does not require 3D detection pretraining.
@@
Comparison on Occ3D-nuScenes without camera mask. To the best of our knowledge, no existing multi-modal occupancy prediction work has used RayIoU metrics to report model performance, so we mainly compare with image-based methods. As shown in Table 2, thanks to the multi-modal input and the effectiveness of our proposed method, our DAOcc is significantly ahead of image-based methods, although many of these image-based methods use multi-frame temporal fusion. It should be noted that some image-based methods based on monocular depth estimation, such as Panoptic-FlashOcc [55], only require images as input during the inference phase, but require additional point cloud data input to provide depth information supervision during the training phase, so these methods can also be regarded as multi-modal during the training phase.
@@
Comparison on SurroundOcc. Table 3 presents a quantitative comparison on SurroundOcc validation set, showcasing the performance of our DAOcc against other methods. By leveraging both LiDAR and camera inputs, DAOcc achieves a remarkable 3.2 mIoU increase over OccFusion [33], which not only employs camera and LiDAR but also incorporates radars. Notably, OccFusion utilizes R101-DCN with a high-resolution input of 900×1600, whereas our method employs a more lightweight Resnet50 backbone and a lower resolution of 256×704. This underscores the effectiveness and efficiency of our proposed method.
@@ 4.4 Ablation Studies
Ablation studies are conducted on Occ3D-nuScenes using the camera mask. Unless otherwise specified, all experiments presented in this section are trained for 15 epochs only.
BEV View Range Voxel Resolution (BEV) mIoU
@@
(b) $\left[-51.2,51.2\right]$ 0.1 $\times$ 0.1 49.98
(c) $\left[-54.0,54.0\right]$ 0.075 $\times$ 0.075 51.26
@@
Table 4: Performance comparison between different BEV processing ranges and voxel sizes.
We propose the BVRE strategy to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective. As shown in Table 4, comparing (a) and (c), increasing the BEV view range from [-41.4, 41.4] to [-54.0, 54.0] can bring a performance improvement of 0.5 mIoU. In addition, by comparing (a) and (b), we can find that the size of the voxel has a great impact on the model’s performance. Setting a smaller voxel size will bring obvious performance improvements but will also impose a greater computational burden.
Detection Auxiliary mIoU bicycle motorcycle traffic cone
51.26 20.60 38.89 36.65
52.82 36.35 45.48 41.98
Table 5: Ablation study on the auxiliary detection head.
To fully leverage the inherent geometric structures within point cloud features, we incorporate 3D object detection as auxiliary supervision. As shown in Table 5, by adding auxiliary supervision for 3D object detection, we can achieve an improvement of 1.56 mIoU, demonstrating the effectiveness of our method. Given that the gain in object detection for foreground objects will be greater, we present the three categories with the highest gain in foreground objects, which are the bicycle (+15.75 mIoU), the motorcycle (+6.59 mIoU) and the traffic cone (+5.33 mIoU). This further proves the rationality of the performance improvement of our method.
Detection Auxiliary EP15 EP24 CBGS (EP6) mIoU

### v2

@@ 3.3 BVRE
As shown in Figure 3, due to the sparsity of point clouds, the increase in computational cost resulting from expanding the processing range of the point cloud is very small compared to increasing the image resolution. Therefore, we extend the point cloud range to provide more 3D spatial context, compensating for the information loss caused by lower image resolution. However, arbitrarily setting the $XY$ range of the point cloud or voxel resolution may result in misalignment between the BEV features and the occupancy ground truth. Assuming that the voxel resolution of the occupancy grid is $res_{o}$, the expanded point cloud data in BEV covers a rectangular region with $X$ ranging from $-x$ to $x$ and $Y$ ranging from $-y$ to $y$, and has a voxel resolution of $res_{p}$. Thus, $x$ and $y$ must be integer multiples of $res_{o}$, and $res_{o}$ must also be an integer multiple of $res_{p}$. Otherwise, the spatial resolution represented by each feature vector along the $z$-axis in the BEV representation will not be equal to $res_{o}$. To avoid complex manual design, we adopt coordinate transformation and interpolation, as illustrated in the purple part of Figure 2, which can be formulated as follows:
$F_{occ}={\rm GridSample}(F_{r},{\rm Norm}(T_{o2l}\times P_{o}))$ (1)
where $P_{o}$ is a set of predefined points in the $XY$ plane of the occupancy annotation coordinate system, each point is located at the center of an occupancy grid in the $XY$ plane. $T_{o2l}$ is the transformation matrix from occupancy coordinates to LiDAR coordinates. The function of ${\rm Norm}$ is to scale the coordinate values to a range from -1 to 1.
3.4 Detachable Auxiliary Detection Head
To fully exploit the inherent geometry and structure information in point cloud features, we employ a detachable auxiliary detection head to further supervise the feature fusion process, thereby enhancing the information about object boundaries and scene structures in the fused features. Meanwhile, the auxiliary detection task is related to the occupancy prediction task, thus providing multiple regularization effects during optimization. For simplicity, we use the one-stage CenterHead introduced in CenterPoint [56] as the auxiliary detection head. Given the refined BEV feature $F_{r}$, we utilize two convolutional layers to generate a keypoint heatmap $HM={p_{xy}}$ from the BEV perspective, and apply a Gaussian kernel [18, 64] to map the center points of all ground truth 3D boxes onto a target heatmap $T$. The training objective is formulated as a focal loss [41] based on Gaussian heatmaps:
@@ 3.4 Detachable Auxiliary Detection Head
$\mathcal{L}_{det}=\mathcal{L}_{cls}+\lambda_{l}\mathcal{L}_{loc}$ (4)
@@ 3.5 Overall Objective Function
Most existing methods rely on complex combinations of loss functions [49, 50, 24, 58, 42, 53, 37, 61, 38, 43], such as a combination of several or all of the focal loss [41], the scene-class affinity loss [4], the dice loss, the lovasz-softmax loss [2], the depth loss [22], etc., to achieve the expected performance. In contrast, our approach utilizes a detachable 3D object detection branch for auxiliary supervision and requires only a simple cross-entropy loss $\mathcal{L}_{ce}$ for occupancy prediction. The total loss for our framework can be defined as:
@@ 4 Experiments
Method mIoU Modality Image Size 2D Backbone
@@
FlashOcc(2f) [57] 43.52 C 512 $\times$ 1408 Swin-B 13.4 51.1 27.7 51.6 56.2 27.3 30.0 29.9 29.8 37.8 43.5 83.8 46.6 56.2 59.6 50.8 44.7
COTR(8f) [36] 46.21 C 512 $\times$ 1408 Swin-B 14.9 53.3 35.2 50.8 57.3 35.4 34.1 33.5 37.1 39.0 45.0 84.5 48.7 57.6 61.1 51.6 46.7
@@
DAOcc (Ours) 53.82 C+L 256 $\times$ 704 R50 12.4 59.6 38.4 61.9 67.1 35.3 48.2 59.1 43.5 50.9 56.3 83.0 44.7 56.7 59.9 70.0 68.1
Table 1: 3D occupancy prediction performance on Occ3D-nuScenes validation set. The camera visible mask is used during the training phase. (xf) means use x frames for temporal fusion. The C, L, and R denote camera, LiDAR, and radar, respectively. Image Size and 2D Backbone represent the input image resolution and the image feature extractor, respectively. The best results are shown in bold.
Method RayIoU Modality Image Size 2D Backbone RayIoU1m RayIoU2m RayIoU4m mIoU
@@
DAOcc (Ours) 48.2 C+L 256 $\times$ 704 R50 44.4 48.7 51.6 47.9
Table 2: 3D occupancy prediction performance on Occ3D-nuScenes validation set. The camera visible mask is not used during the training phase. (xf) means use x frames for temporal fusion. The C and L denote camera, LiDAR, respectively. The best results are shown in bold.
Method IoU mIoU Modality Image Size 2D Backbone
@@
DAOcc (Ours) 45.0 30.5 C+L 256$\times$ 704 R50 30.8 19.5 34.0 38.8 25.0 27.7 29.9 22.5 23.2 31.6 41.0 25.9 29.4 29.9 35.2 43.5
Table 3: 3D occupancy prediction performance on SurroundOcc validation set. The C, L, and R denote camera, LiDAR, and radar, respectively. The best results are shown in bold.
@@ 4.1 Datasets and Evaluation Metrics
Occ3D-nuScenes [44] is a large-scale benchmark dataset for 3D occupancy prediction, which contains 700 scenes for training and 150 scenes for validation. Each scene lasts for 20 seconds, with annotations provided at a frequency of 2 Hz. The dataset covers a perception range of [-40m, 40m] in the $X$ and $Y$ directions, and [-1m, 5.4m] along the $z$-axis, all within the ego coordinate system. This range is discretized into voxels, each with a size of [0.4m, 0.4m, 0.4m]. Each occupied voxel is assigned one of 17 semantic labels, including 16 common classes and a general object class labeled as ‘others’. The dataset includes RGB images captured by six surrounding cameras and LiDAR point clouds, enabling dense voxel-wise annotations for 3D scene understanding. Additionally, the dataset provides visibility masks for both LiDAR and camera modes, which can be utilized for training purposes.
SurroundOcc [51] is an automatically generated occupancy prediction dataset that does not require extra human annotation. It is built upon the nuScenes [3] and generates dense occupancy ground truth by utilizing existing 3D detection and 3D semantic segmentation labels. The dataset uses the nuScenes training set to train the model and employs the nuScenes validation set for evaluation purposes. The occupancy prediction range is set to [-50m, 50m] for the $X$ and $Y$ axes, and [-5m, 3m] for the $z$-axis, all within the LiDAR coordinate system. The final output 3D occupancy grid has a shape of 200×200×16, with each voxel measuring [0.5m, 0.5m, 0.5m].
Evaluation metrics: Following previous methods, we use the mean Intersection over Union (mIoU) as our evaluation metric when training models with camera visible mask provided by Occ3D-nuScenes, ensuring a fair comparison with existing approaches. When the camera visible mask is excluded from the training phase, we report an additional semantic segmentation metric, RayIoU [31], which has been recently introduced. For the SurroundOcc dataset, we report semantic segmentation performance using IoU and mIoU. Here, IoU is defined as the intersection over union of occupied voxels, calculated while ignoring their semantic categories, and it is used as an evaluation metric for the scene completion (SC) task.
@@ 4.2 Implementation Details
We use ResNet-50 [12], pre-trained on nuImages [3] by MMDetection [5], as our image backbone and adjust the input image resolution to 256×704. In the LiDAR branch, we aggregate LiDAR sweeps from 10 adjacent timestamps as input and initiate training from scratch. Data augmentation techniques, such as scaling, rotation, and flipping, are applied to the input images, and we also perform random flipping in the BEV space. The loss weights are configured as $\lambda_{l}=0.25$ and $\lambda=0.01$. We adopt the AdamW optimizer [35] with a cosine annealing learning rate scheduler, incorporating a warmup phase, and set the initial learning rate to 2e-4. Unless conducting ablation experiments, we train our DAOcc using CBGS [65] for 6 epochs, the input point cloud range is set to [-54.0, -54.0, -5.0, 54.0, 54.0, 3.0] and the voxel size is set to [0.075, 0.075, 0.2]. All models are trained on 8 NVIDIA RTX4090 GPUs with a batch size of 4. Additionally, if the camera visible mask is not used during training, we randomly discard 80% of empty voxels, consistent with Zhao et al. [63]. Temporal information fusion, exponential moving average (EMA) for weight parameter and test-time augmentation is not used.
@@ 4.3 Comparison with State-of-the-Art Methods
Method Modality 3D Backbone Param mIoU
FB-Occ [24] C - 130.8M 48.90
RadOcc [60] C+L SECOND 135.4M 49.38
EFFOcc [43] C+L SECOND 108.6M 52.62
DAOcc C+L SECOND 56.2M 53.82
Table 4: Complexity comparisons. The C and L denote camera, LiDAR, respectively. 3D Backbone represents the network used for point cloud feature extraction. The reported performance mIoU is the result of training with camera mask on Occ3D-nuScenes validation set.
@@
Comparison on Occ3D-nuScenes with camera mask. As presented in Table 1, we compare our proposed DAOcc with both image-based and multi-modal methods on the Occ3D-nuScenes dataset. All results are provided either directly by their authors or based on their official code implementations. Our DAOcc establishes a new state-of-the-art performance, achieving an outstanding 53.82 mIoU, while utilizing a 2D backbone ResNet50 [12] and 256×704 input image resolution. Compared with the latest state-of-the-art multi-modal method EFFOcc [43], DAOcc achieves improvement of 1.2 mIoU. Notably, the image feature extraction network Swin-B [32] in EFFOcc has approximately 3 times the number of parameters of ResNet50, and EFFOcc uses a larger input image resolution of 512×1408. This highlights our method’s ability to mine more useful information from point cloud data. In addition, to achieve optimal results, EFFOcc requires the LiDAR branch to undergo 3D object detection pre-training, which leads to an improvement of approximately 3 mIoU. In contrast, DAOcc does not require LiDAR branch pre-training, but instead seamlessly integrates 3D object detection as an auxiliary branch into the multi-modal occupancy prediction framework, simplifying the model training process and achieving better performance.
Furthermore, we compare the complexity of the models, as shown in Table 4. Compared to the two top-performing multi-modal methods, RadOcc [60] and EFFOcc [43], as well as the top-performing image-based method FB-Occ [24], our proposed DAOcc has only about half the number of parameters but achieves significantly better performance (outperforming RadOcc, EFFOcc, and FB-Occ by 4.44, 1.2 and 4.92 mIoU).
@@
Comparison on Occ3D-nuScenes without camera mask. To the best of our knowledge, no existing multi-modal occupancy prediction work has used RayIoU metrics to report model performance, so we mainly compare with image-based methods. As shown in Table 2, benefiting from the effective use of multi-modal input, especially point cloud data, our DAOcc achieves an outstanding 48.2 RayIoU and significantly outperforms image-based methods by a large margin, even though many of these methods use multi-frame temporal fusion. It should be noted that some image-based methods, such as Panoptic-FlashOcc and FB-Occ [24], adopt the Explicit Depth Supervision [22] and leverage depth information derived from point cloud data to supervise the training of monocular depth estimation from images, so these methods can also be regarded as multi-modal during the training phase.
@@
Comparison on SurroundOcc. Table 3 presents a quantitative comparison on the SurroundOcc validation set, showcasing the performance of our DAOcc against other methods. Leveraging both camera and LiDAR inputs, DAOcc achieves a notable 3.2 mIoU improvement over OccFusion [37], which incorporates radar data in addition to camera and LiDAR inputs. It is worth noting that OccFusion utilizes R101-DCN with a high-resolution input of 900×1600, whereas our method employs a more lightweight Resnet50 backbone and a lower input resolution of 256×704. This underscores the effectiveness and efficiency of our proposed method.
@@ 4.4 Ablation Studies
Ablation studies are conducted on the validation set of Occ3D-nuScenes, with the camera visible mask being used during training. Unless otherwise specified, all experiments presented in this section are trained for 15 epochs without employing CBGS [65].
BEV View Range (m) BEV Voxel Size (m) mIoU
@@
(b) $\left[-45.6,45.6\right]$ 0.075 $\times$ 0.075 50.98
(c) $\left[-45.6,45.6\right]$ 0.1 $\times$ 0.1 49.75
(d) $\left[-51.2,51.2\right]$ 0.1 $\times$ 0.1 49.98
(e) $\left[-54.0,54.0\right]$ 0.075 $\times$ 0.075 51.26
@@
Table 5: Performance comparisons between different BEV ranges and voxel sizes. In all experiments, the voxel height is 0.2m.
Effectiveness of the BVRE. We propose the BVRE strategy to enrich spatial contextual information by expanding the perceptual range from a BEV (Bird’s Eye View) perspective. As shown in Table 5, a comparison of (a) and (b) indicates that, with the same voxel size, expanding the BEV range from [-41.4m, 41.4m] to [-45.6m, 45.6m] yields a performance improvement of 0.22 mIoU. Furthermore, the comparison of (a) and (e) shows that further expanding the BEV range to [-54.0m, 54.0m] augments the performance gain to 0.5 mIoU. This observation validates the value of spatial context information in a larger BEV range.
In addition, by comparing (b) and (c), we can find that the size of the voxel has a great impact on the model’s performance. To achieve better performance, we adopt the setting (e) in DAOcc.
camera mask ADH mIoU
$\blacksquare$ bicycle
$\blacksquare$ motorcycle
$\blacksquare$ pedestrian
$\blacksquare$ traffic cone
51.26 20.60 38.89 55.44 36.65
52.82 (+1.56) 36.35 (+15.75) 45.48 (+6.59) 58.27 (+2.83) 41.98 (+5.33)
44.31 14.94 32.22 43.77 34.07
46.89 (+2.58) 32.51 (+17.57) 42.89 (+10.67) 53.41 (+9.64) 40.55 (+6.48)
Table 6: Ablation study on the Auxiliary Detection Head. ADH is short for the Auxiliary Detection Head.
Ablation on the Auxiliary Detection Head. To fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision. As shown in Table 6, by adding auxiliary 3D object detection supervision, we can achieve an improvement of 1.56 mIoU when using camera visible masks during training and 2.58 mIoU when not using the mask, respectively. This demonstrates that adding a detection branch can significantly enhance the discriminability of the multi-modal features. Furthermore, we report on the four categories that benefit the most from the 3D object detection supervision. Notably, all four categories are foreground objects, which further validates the effectiveness and rationale behind the performance improvement achieved by our method. Specifically, when the camera visible mask is used during training, the addition of the auxiliary detection branch results in improvements of 15.75 mIoU, 6.59 mIoU, 2.83 mIoU, and 5.33 mIoU for bicycles, motorcycles, pedestrians, and traffic cones, respectively. Conversely, when the camera visible mask is not used during training, the improvements are 17.57 mIoU, 10.67 mIoU, 9.64 mIoU, and 6.48 mIoU for the same categories, respectively.
5 Conclusion
In this paper, we propose a novel multi-modal occupancy prediction framework, DAOcc, which aims to achieve superior performance while utilizing a deployment-friendly image backbone and practical input image resolutions. DAOcc achieves this superior performance by fully leveraging the unique advantages of point cloud data within a multi-modal framework, advantages that have been largely overlooked by previous methods. Extensive experiments conducted on the Occ3D-nuScenes and SurroundOcc datasets demonstrate the superiority of DAOcc compared to existing methods. We believe that DAOcc has practical application potential and could inspire future research in multi-modal occupancy prediction.
Supplementary Material
Method Modality GPU Type FPS mIoU
FlashOcc(2f) [57] C A100-PCIE-40G 0.9 43.52
RadOcc [60] C+L - 0.3 49.38
EFFOcc [43] C+L RTX A6000 1.8 52.62
DAOcc(ours) C+L RTX 3070Ti 2.7 53.82
Table 7: Comparisons of the runtime. (xf) means use x frames for temporal fusion. The C and L denote camera, LiDAR, respectively. FPS for DAOcc is measured with the PyTorch fp32 backend. The reported performance mIoU is the result of training with camera mask on Occ3D-nuScenes validation set.
6 Runtime Comparisons
We compare the runtime of different methods, as shown in Table 7. The FPS is obtained by testing on the corresponding GPU Type listed in the table. Since the source code for RadOcc is unavailable and EFFOcc has not yet released its implementation, we rely on the runtime speeds reported in their respective papers as benchmarks. It’s worth noting that the computing capability of the RTX 3070 Ti is significantly lower than that of both the RTX A6000 and the A100. The results demonstrate that our proposed DAOcc achieves the best performance while maintaining the fastest runtime speed.
7 Ablation for Achieving Final Performance
In addition to the ablation studies presented in Section 4.4, this section provides further ablation experiments. Table 8 presents the ablation experiments that contributed to the final performance. First, by extending the training time from 15 epochs to 24 epochs, we can observe an improvement of 0.57 mIoU. Next, the use of CBGS [65] yields an additional 0.43 mIoU improvement. Since CBGS involves resampling, which increases the number of samples per epoch, we limit the training to 6 epochs to maintain a comparable total training time to the 24 epochs without CBGS for a fair comparison.
8 Detection Performance
Since our proposed DAOcc incorporates an auxiliary 3D object detection branch into the occupancy prediction framework, we also evaluate the performance of this auxiliary branch. The auxiliary detection branch is designed to identify 10 object categories, including cars, trucks, buses, trailers, construction vehicles, pedestrians, motorcycles, bicycles, traffic cones, and barriers. All experiments are conducted using the nuScenes [3] dataset for the auxiliary 3D object detection training, and the results are evaluated using the mean average precision (mAP) and nuScenes detection score (NDS) metrics provided by the nuScenes benchmark.
As shown in Table 9, when DAOcc is trained on the Occ3D-nuScenes [44] dataset (for occupancy prediction) and the camera visible mask is used during training, the auxiliary detection branch achieves a performance of 59.4 mAP and 64.3 NDS. Notably, comparable performance is observed when the camera visible mask is not used. However, when the SurroundOcc [51] dataset is employed for occupancy prediction training, the auxiliary detection branch achieves a lower performance of 53.6 mAP and 60.2 NDS. We attribute this performance gap to the fact that the surroundocc dataset is automatically generated without manual calibration, potentially introducing more noise and negatively affecting model convergence.
EP15 EP24 CBGS (EP6) mIoU